the sides of a triangle are 3x+4y, 4x+3y, 5x+5y where x,y>0. then the triangle is 
  1. right angled
  2. obtude angled
  3. equilateral
  4. none of these.

Grade:11

5 Answers

Ravi
9 years ago
y=-3x/4 slope= -3/4
y=-4x/3 slope=-4/3
y=-x slope=-1

Find the angle between two sides using the formula
tan A= (m1-m2)/(1+m1m2)

This will give angle A between the 2 sides whose slope has been taken in the formula as m1,2.


Find for any 2 angles and find the third angle using the Angle sum property and hence determine the type of triangle.


Hint for shortcut: All the lines are passing through a common point (0,0)
Ravi
9 years ago
y=-3x/4 slope= -3/4
y=-4x/3 slope=-4/3
y=-x slope=-1

Find the angle between two sides using the formula
tan A= (m1-m2)/(1+m1m2)

This will give angle A between the 2 sides whose slope has been taken in the formula as m1,2.


Find for any 2 angles and find the third angle using the Angle sum property and hence determine the type of triangle.


Hint for shortcut: All the lines are passing through a common point (0,0)
Ravi
9 years ago
y=-3x/4 slope= -3/4
y=-4x/3 slope=-4/3
y=-x slope=-1

Find the angle between two sides using the formula
tan A= (m1-m2)/(1+m1m2)

This will give angle A between the 2 sides whose slope has been taken in the formula as m1,2.


Find for any 2 angles and find the third angle using the Angle sum property and hence determine the type of triangle.


Hint for shortcut: All the lines are passing through a common point (0,0)
Ravi
9 years ago
y=-3x/4 slope= -3/4
y=-4x/3 slope=-4/3
y=-x slope=-1

Find the angle between two sides using the formula
tan A= (m1-m2)/(1+m1m2)

This will give angle A between the 2 sides whose slope has been taken in the formula as m1,2.


Find for any 2 angles and find the third angle using the Angle sum property and hence determine the type of triangle.


Hint for shortcut: All the lines are passing through a common point (0,0)
sajid
8 years ago
it is a right angled triangle